# Jaipur-302017, INDIA

### v2

@@ 3 Ungapped phases
4 Gapped Phase
Figure 3: Schematic of the support $\mathcal{C}$ of the spectral function in the gapped phase (dashed, both panels). The dots on the negative real axis are the poles of the potential ($z=-1/\alpha,-1/\gamma$). Solid closed curve in the left panel is the contour $\mathcal{C}^{*}$ used to solve integrals in the gapped phase. It can be deformed continuously to that shown in the right panel. See Fig. 4 for the actual contours in the gapped phase.
@@ 4 Gapped Phase
$R(z)=\left\langle\frac{z+U}{z-U}\right\rangle.$ (30)
@@
$R(z)=\int_{\mathcal{C}}\frac{dz^{\prime}}{2\pi iz^{\prime}}\,\rho(z^{\prime})\,\frac{z+z^{\prime}}{z-z^{\prime}}.$ (32)
@@
Figure 4: Support of the spectral function $\rho(z)$ in the gapped phase.
By using the Sokhotski-Plemelj formula Muskhelishvili (2013), we find the two main expressions which make the resolvent a useful quantity. The first relates the potential to the resolvent:
$zV^{\prime}(z)=\frac{1}{2}\,\lim_{\epsilon\to 0}\Big(R(z+\epsilon)+R(z-\epsilon)\Big).$ (33)
@@
$\rho(z)=\frac{1}{2}\Delta_{\mathcal{C}}R(z)$ (34)
@@
$R(z)=zV^{\prime}(z)+g(z)Q(z).$ (35)
Note that $V$ in the above is the modified potential which includes the Lagrange multplier, as written in (4). The second term contains $g(z)$, a unknown function well-behaved on $\mathcal{C}$ but possibly meromorphic outside it, and $Q(z)$, which contains the discontinuity in $R(z)$. The latter is defined as $Q(z)=\sqrt{z-a}\,\sqrt{z-a^{*}}$ with branch cut $\mathcal{C}$. This definition ensures that $\rho(a)=\rho(a^{*})=0$. Using the ansatz in (34) gives
$\rho(z)=\frac{1}{2}g(z)\,\Delta_{\mathcal{C}}Q(z).$ (36)
@@
To find an expression for $g(z)$, we first note that (34) allows writing the expectation value of a quantity $f(U)$ as a closed contour integral:
$\langle f(U)\rangle=\int_{\mathcal{C}}\dfrac{dz}{2\pi iz}\,\rho(z)\,f(z)=\oint_{\mathcal{C}^{*}}\dfrac{dz}{4\pi iz}\,R(z)\,f(z),$ (37)
@@
$g(z)=\oint_{\mathcal{C}^{*}}\frac{z^{\prime}V(z^{\prime})}{(z^{\prime}-z)Q(z^{\prime})}\frac{dz^{\prime}}{2\pi i}.$ (38)
This result is similar to a theorem by Tricomi G. Tricomi (1985).
Figure 5: Phase diagram of the model for $\sigma=0.5$ and $\sigma=2$. The four ungapped phases are represented by various colors and the gapped phase by white. As $\sigma$ is increased the gapped phase grows outwards from the center at $\alpha=1,\gamma=1$.
Figure 6: Wilson loop (left panels) and Lagrange multiplier (right panels) as the system transitions from ungapped phase A to B (top panels) and from phase D to C (bottom panels) through the gapped phase.
For the current model, $\mathcal{C}^{*}$ can be continuously deformed to that shown in the right panel, which means that the integral in the above relation takes contributions only from the residues at $z=0,-\frac{1}{\alpha},-\frac{1}{\gamma},\infty$. Solving the integral we find
$g(z)=\frac{\sigma}{(1+\gamma z)A_{\gamma}}+\frac{\sigma}{(1+\alpha z)A_{\alpha}}$ (39)
@@
$A_{\alpha}=\left|\frac{1}{\alpha}+a\right|,\quad A_{\gamma}=\left|\frac{1}{\gamma}+a\right|.$
@@
$\mathscr{N}+\sigma-\sigma|a|\left(\frac{1}{A_{\alpha}}+\frac{1}{A_{\gamma}}\right)=-1,\quad\mathscr{N}-\sigma+\sigma\left(\frac{1}{\alpha\,A_{\alpha}}+\frac{1}{\gamma A_{\gamma}}\right)=1.$ (40)
The third and last equation is found by enforcing the SU($N$) constraint $\langle\ln U\rangle=0$. By using (37) this constraint can be rewritten in the following form
$\oint_{\mathcal{C}^{*}}\dfrac{dz}{4\pi iz}\,R(z)\,\ln z=0$ (41)
@@
$\dfrac{1}{A_{\alpha}}\left[\left(|a|+\dfrac{1}{\alpha}\right)\ln|a|-\left(|a|-\dfrac{1}{\alpha}\right)\ln\cos^{2}\frac{\phi}{2}\right]-\text{arcsinh}\dfrac{(\alpha-|a|^{-1})A_{\alpha}}{1+\cos\phi}+\ln\alpha$ (42)
$+\dfrac{1}{A_{\gamma}}\left[\left(|a|+\dfrac{1}{\gamma}\right)\ln|a|-\left(|a|-\dfrac{1}{\gamma}\right)\ln\cos^{2}\frac{\phi}{2}\right]-\text{arcsinh}\dfrac{(\gamma-|a|^{-1})A_{\gamma}}{1+\cos\phi}+\ln\gamma=0,$
@@
$W_{1}=\dfrac{\sigma}{2}\left(\dfrac{1}{\alpha}+\dfrac{1}{\gamma}\right)-\dfrac{\sigma}{2|a\alpha+1|}\left(\dfrac{1}{\alpha}+\text{Re}\,a\right)-\dfrac{\sigma}{2|a\gamma+1|}\left(\dfrac{1}{\gamma}+\text{Re}\,a\right)$ (43)
$W_{-1}=\dfrac{\sigma}{2}(\alpha+\gamma)-\dfrac{\sigma\alpha}{2|\alpha+1/a|}\left(\alpha+\dfrac{\text{Re}\,a}{|a|^{2}}\right)-\dfrac{\sigma\gamma}{2|\gamma+1/a|}\left(\gamma+\dfrac{\text{Re}\,a}{|a|^{2}}\right)$ (44)
@@
$\frac{\partial F}{\partial\alpha}=-\frac{\sigma^{2}}{2\alpha}\left(\frac{1}{\alpha A_{\alpha}}+\frac{1}{\gamma A_{\gamma}}\right)+\frac{\sigma^{2}}{2(\alpha-\gamma)}\left(1-\frac{A_{\gamma}}{A_{\alpha}}\right)+\frac{\sigma^{2}}{2\alpha^{2}A^{2}_{\alpha}}\left(\frac{1}{\alpha}+\text{Re}\,a\right).$ (45)
The support, found by solving (14), is shown in Fig. 4. All plots for the gapped phase were made using Mathematica software. We discuss the phase transitions in the model next.
5 Phase transitions
@@
Figure 7: The beta function (top panels) and the first two derivatives of the free energy (bottom four panels) are continuous at the transition point. However, the beta function and the second derivative of the free energy show kinks at the transition points (the inset in the bottom panels zooms in on the kink). These kinks signal a third order transition.
@@
This model, similar to the Gross-Witten-Wadia model, exhibits third order phase transitions between the ungapped and gapped phases. Fig. 5 shows the full phase diagram. Interestingly, there are no direct transitions between the ungapped phases. We also note that as the coupling $\sigma$ increases, the gapped phase expands outwards from the center, while the ungapped phases shrink. At $\sigma=0$ a single ungapped phase exists, while at $\sigma=\infty$ only the gapped phase remains. All five phases exist at every finite value of $\sigma$. Fig. 6 shows the Wilson loop expectation values and the Lagrange multiplier as the system transitions from one ungapped phase to another through the intermediate gapped phase.
@@
The first derivative of the beta function, the second derivative of the Lagrange multiplier $\mathscr{N}$ and the third derivative of the free energy are all discontinuous at the transition, indicating that the phase transition is of order three. Fig. 7 shows the beta function and the derivatives of $F$. Although the plot shows derivatives of $F$ with respect to only $\alpha$, the discontinuity in the third derivative $F$ exists for all parameters, including combinations such as $\frac{\partial^{3}F}{\partial\alpha\,\partial^{2}\gamma}$.
Lastly, as mentioned in the introduction, our model converges to the low temperature two-level QCD model of Ref. Hands et al. (2010) if we relate the parameters $\alpha,\gamma$ to the temperature $T$ and the chemical potential $\mu$ as follows
$\alpha=e^{(\mu-\varepsilon)/T},\qquad\gamma=e^{(\mu+\varepsilon)/T}$ (46)
where $\varepsilon$ is the energy difference between the two levels. In this case, $\sigma$ is proportional to the number of quark flavors, $\mathscr{N}$ becomes the fermion number, and the ungapped and gapped phases have the interpretation of confined and deconfined phases. In the exact zero temperature limit $\gamma$ approaches infinity, and one of the log terms in the potential (3) can be dropped, which is exactly what the authors of Hands et al. (2010) have done. Keeping both terms produces the phase diagram shown in Fig. 8, which differs from that of Hands et al. (2010) only at large temperatures. Only two confined phases exist in this diagram, as the other two correspond to negative temperature. We also note that as $\sigma$ increases, the deconfined phase expands, while the confined phase shrinks to smaller values of $T$. The three phases exist at every finite value of $\sigma$. Lastly, it appears that a direct confined to confined transition is allowed at $T=0$. This apparent contradiction with our earlier discussion is explained by the fact that the potential (3) does not really cover the $T=0$ case as the parameters $\alpha$ and $\gamma$ are ill-defined at $T=0$.
@@
Figure 8: $T$-$\mu$ phase diagram of the model when related to QCD. There are two confined phases (colored) and one deconfined phase.
6 Conclusions and Discussions
@@
In this work, we have examined the large $N$ limit of a SU($N$) unitary matrix model with a complex potential containing Fisher-Hartwig singularities. The model exhibits four ungapped phases and a single gapped phase. The saddle points in the large $N$ limit were found to lie off the unit circle in the complex plane for both ungapped and gapped phases, reflecting the inherent complex-valued nature of the model’s action. Transitions between the ungapped phases proceed through the intermediate gapped phase. The phase transitions are third order, similar to that observed in the Gross–Witten-Wadia model.
@@
Exact and explicit expressions were obtained for all observables in the ungapped phases, whereas in the gapped phase the transcendental nature of the equations forced us to use numerical methods to study the observables. We have also connected our results to an existing low temperature QCD-like model to verify our results, and extend the existing results.
@@
These results provide deeper insight into the phase structure of complex extensions of unitary matrix models with Fisher-Hartwig singularities and may offer useful analogies for understanding complex actions in QCD-like theories. Further extensions of this work could include incorporating finite-$N$ effects or exploring the behavior under further complex deformations of the potential.

### v3

@@ 3.1. Ungapped phases
3.2. Gapped Phase
Figure 5. Support of the spectral function $\rho(z)$ in the gapped phase.
@@ 3.2. Gapped Phase
(41) $R(z)=\left\langle\frac{z+U}{z-U}\right\rangle.$
@@
(43) $R(z)=\int_{\mathcal{C}}\frac{dz^{\prime}}{2\pi iz^{\prime}}\,\rho(z^{\prime})\,\frac{z+z^{\prime}}{z-z^{\prime}}.$
@@
Figure 6. Large-NN phase diagram of the model for $\sigma=0.5$ and $\sigma=2$. The four ungapped phases are represented by various colors and the gapped phase by white. As $\sigma$ is increased the gapped phase grows outwards from the center $\alpha=1,\gamma=1$.
By using the Sokhotski-Plemelj formula [29], we find the two main expressions which make the resolvent a useful quantity. The first relates the potential to the resolvent:
(44) $zV^{\prime}(z)=\frac{1}{2}\,\lim_{\epsilon\to 0}\Big(R(z+\epsilon)+R(z-\epsilon)\Big).$
@@
(45) $\rho(z)=\frac{1}{2}\Delta_{\mathcal{C}}R(z)$
@@
(46) $R(z)=zV^{\prime}(z)+g(z)Q(z).$
The second term contains $g(z)$, an unknown function well-behaved on $\mathcal{C}$ but possibly meromorphic outside it, and $Q(z)$, which contains the discontinuity in $R(z)$. The latter is defined as $Q(z)=\sqrt{z-a}\,\sqrt{z-a^{*}}$ with branch cut $\mathcal{C}$. This definition ensures that $\rho(a)=\rho(a^{*})=0$. Using the ansatz in (45) gives
(47) $\rho(z)=\frac{1}{2}g(z)\,\Delta_{\mathcal{C}}Q(z).$
@@
Figure 7. Wilson loop (left panels) and Lagrange multiplier (right panels) in the large-$N$ limit as the system transitions from ungapped phase A to B (top panels) and from phase D to C (bottom panels) through the gapped phase.
To find an expression for $g(z)$, we first note that (45) allows writing the expectation value of a quantity as a closed contour integral:
(48) $\frac{1}{N}\langle\mathrm{Tr}\,h(U)\rangle=\int_{\mathcal{C}}\dfrac{dz}{2\pi iz}\,\rho(z)\,h(z)=\oint_{\mathcal{C}^{*}}\dfrac{dz}{4\pi iz}\,R(z)\,h(z),$
@@
(49) $g(z)=\oint_{\mathcal{C}^{*}}\frac{z^{\prime}V(z^{\prime})}{(z^{\prime}-z)Q(z^{\prime})}\frac{dz^{\prime}}{2\pi i}.$
This result is similar to a theorem by Tricomi [16]. For the current model, $\mathcal{C}^{*}$ can be continuously deformed to that shown in the right panel, which means that the integral in the above relation takes contributions only from the residues at $z=0,-\frac{1}{\alpha},-\frac{1}{\gamma},\infty$. Solving the integral we find
(50) $g(z)=\frac{\sigma}{(1+\gamma z)B_{\gamma}}+\frac{\sigma}{(1+\alpha z)B_{\alpha}}$
@@
$B_{\alpha}=\left|\frac{1}{\alpha}+a\right|,\quad B_{\gamma}=\left|\frac{1}{\gamma}+a\right|.$
@@
(51) $\mathscr{N}+\sigma-\sigma|a|\left(\frac{1}{B_{\alpha}}+\frac{1}{B_{\gamma}}\right)=-1,\quad\mathscr{N}-\sigma+\sigma\left(\frac{1}{\alpha\,B_{\alpha}}+\frac{1}{\gamma B_{\gamma}}\right)=1.$
The third and last equation is found by enforcing the SU($N$) constraint $\langle\mathrm{Tr}\,\ln U\rangle=0$. By using (48) this constraint can be rewritten in the following form
(52) $\oint_{\mathcal{C}^{*}}\dfrac{dz}{4\pi iz}\,R(z)\,\ln z=0$
@@
(53) $\dfrac{1}{B_{\alpha}}\left[\left(|a|+\dfrac{1}{\alpha}\right)\ln|a|-\left(|a|-\dfrac{1}{\alpha}\right)\ln\cos^{2}\frac{\phi}{2}\right]-\text{arcsinh}\dfrac{(\alpha-|a|^{-1})B_{\alpha}}{1+\cos\phi}+\ln\alpha$
$+\dfrac{1}{B_{\gamma}}\left[\left(|a|+\dfrac{1}{\gamma}\right)\ln|a|-\left(|a|-\dfrac{1}{\gamma}\right)\ln\cos^{2}\frac{\phi}{2}\right]-\text{arcsinh}\dfrac{(\gamma-|a|^{-1})B_{\gamma}}{1+\cos\phi}+\ln\gamma=0,$
@@
(54) $W_{1}=\dfrac{\sigma}{2}\left(\dfrac{1}{\alpha}+\dfrac{1}{\gamma}\right)-\dfrac{\sigma}{2|a\alpha+1|}\left(\dfrac{1}{\alpha}+\text{Re}\,a\right)-\dfrac{\sigma}{2|a\gamma+1|}\left(\dfrac{1}{\gamma}+\text{Re}\,a\right)$
(55) $W_{-1}=\dfrac{\sigma}{2}(\alpha+\gamma)-\dfrac{\sigma\alpha}{2|\alpha+1/a|}\left(\alpha+\dfrac{\text{Re}\,a}{|a|^{2}}\right)-\dfrac{\sigma\gamma}{2|\gamma+1/a|}\left(\gamma+\dfrac{\text{Re}\,a}{|a|^{2}}\right)$
@@
(56) $\frac{\partial F}{\partial\alpha}=-\frac{\sigma^{2}}{2\alpha}\left(\frac{1}{\alpha B_{\alpha}}+\frac{1}{\gamma B_{\gamma}}\right)+\frac{\sigma^{2}}{2(\alpha-\gamma)}\left(1-\frac{B_{\gamma}}{B_{\alpha}}\right)+\frac{\sigma^{2}}{2\alpha^{2}A^{2}_{\alpha}}\left(\frac{1}{\alpha}+\text{Re}\,a\right).$
The support, found by numerically solving (27), is shown in Fig. 5. All plots for the gapped phase were made using Mathematica software. In the next section, we discuss the phase transitions in the model.
@@
Figure 8. The first two derivatives of the large-$N$ free energy are continuous at the transition point. However, the second derivative of the free energy has a kink at the transition points, which signals a third order transition. The inset in the bottom panels zooms in on the kink.
4. Summary of phase transitions
@@
This model, similar to the Gross-Witten-Wadia model, exhibits third order phase transitions in the large-$N$ limit. See Fig. 6 for the full large-$N$ phase diagram. The present model addtionally exhibits finite-$N$ phase transitions whenever the inverse of the coupling is a non-integer, due to the non-analyticity of the potential at the branch cuts. The order of these finite-$N$ transitions depends on the coupling $g$ and is given by $1+\mathrm{ceiling}(1/g)$. This fact ensures that the transitions disappear in the double scaling large-$N$ limit, and they are replaced by an ungapped-to-gapped transition. Interestingly, there are no direct transitions between the different ungapped phases. We also note that as the coupling $\sigma$ increases, the gapped phase expands outwards from the center, while the ungapped phases shrink. At $\sigma=0$ a single ungapped phase exists, while at $\sigma=\infty$ only the gapped phase remains. All five phases exist at every finite value of $\sigma$. Fig. 7 shows the Wilson loop expectation values and the Lagrange multiplier as the system transitions from one ungapped phase to another through the intermediate gapped phase. The Wilson loop $W_{1}$ in particular exhibits a behaviour similar to the famous Silver Blaze phenomenon of QCD, where a thermodynamic quantity is independent of a parameter (the chemical potential to be precise) on one side of the transition, and dependent on the other, despite the fact that the parameter appears explicitly in the action.
@@
In the large-$N$ limit, the second derivative of the Lagrange multiplier $\mathscr{N}$ and the third derivative of the free energy are discontinuous at the transition, indicating that the phase transition is third order. See Fig. 8 for plots of the derivatives of $F$. Although derivatives of $F$ with respect to only $\alpha$ are shown, the discontinuity in the third derivative $F$ exists for all parameters, including combinations such as $\frac{\partial^{3}F}{\partial\alpha\,\partial^{2}\gamma}$.
@@
Figure 9. $T$-$\mu$ phase diagram of the large-$N$ model when related to weakly-coupled QCD. There are two confined phases (colored) and one deconfined phase.
Finally, as mentioned in the introduction, our model converges to the low temperature two-level QCD model of Ref. [21] if we relate the parameters $\alpha,\gamma$ to the temperature $T$ and the chemical potential $\mu$ as follows
(57) $\alpha=e^{(\mu-\varepsilon)/T},\qquad\gamma=e^{(\mu+\varepsilon)/T}$
where $\varepsilon$ is the energy difference between the two levels. In this case, $\sigma$ is proportional to the number of quark flavors, $\mathscr{N}$ becomes the fermion number. In this interpretation, the finite-$N$ transitions occur at $\mu_{c}=\varepsilon$. The large-$N$ ungapped and gapped phases have the interpretation of confined and deconfined phases. In the exact zero-temperature limit $\gamma$ approaches infinity, and one of the log terms in the potential (17) can be dropped, which is exactly what the authors of Ref. [21] have done. Keeping both terms produces the phase diagram shown in Fig. 9, which agrees with that in Ref. [21] at small temperatures. Only two confined phases exist in this diagram, as the other two correspond to negative temperature. We also note that as $\sigma$ increases, the deconfined phase expands, while the confined phase shrinks to smaller values of $T$. The three phases exist at every finite value of $\sigma$. Lastly, it appears in the plots that a direct confined to confined transition is allowed at $T=0$. This apparent contradiction with the previous discussion is resolved by noting that the potential (17) does not really cover the $T=0$ case. Specifically, the mapping to the parameters $\alpha$ and $\gamma$ given in (57) is pathological at $T=0$.
5. Conclusion
@@
In this work, we have examined an SU($N$) unitary matrix model with a complex action containing Fisher-Hartwig singularities. The model exhibits phase transitions of coupling-dependent order at finite-$N$. These transitions disappear in the exact large-$N$ limit, and replaced by third order Gross-Witten-Wadia transitions between four ungapped phases and a single gapped phase. Transitions between the ungapped phases proceed through the intermediate gapped phase, never directly. The saddle points in the large-$N$ limit were found to lie off the unit circle in the complex plane for both ungapped and gapped phases, reflecting the inherent complex-valued nature of the model’s action.
@@
Exact and explicit expressions were obtained for all observables in the ungapped phases, whereas in the gapped phase the transcendental nature of the equations forced us to use numerical methods to study the observables. We have also connected our results to an existing model of low-temperature weak-coupling QCD to verify our results, and extended the existing results.
@@
These results provide deeper insight into the phase structure of complex extensions of unitary matrix models with Fisher-Hartwig singularities and may offer useful analogies for understanding complex actions in QCD-like theories. Further extensions of this work could include complete expressions for the finite-$N$ observables and their detailed anaylsis, or exploring the behavior under further complex deformations of the potential.